OpenStudy (anonymous):

During the Scouts Week, Mike washed cars and mowed the lawn. He charged $9 for each car washed and $5 for each lawn mowed. (a)How many cars did he wash if he earned $252 for washing cars only? (B) If he earned $387 altogether, how many lawns did he mow?

OpenStudy (hihi67):

Okay for (a) you have to divide 252 by 9 (the price of a car wash.)

OpenStudy (hihi67):

Do you get it @t0304427g

OpenStudy (anonymous):

ok

OpenStudy (hihi67):

??? @elyn

OpenStudy (anonymous):

how about the (B)

OpenStudy (hihi67):

Well, to find out how much he earned mowing lawns, you have to subtract 387 with 252.

OpenStudy (hihi67):

After you have done that, you can do the same as in (a) which is to divide the price of mowing lawns (5) by your new answer.

OpenStudy (anonymous):

A= 252/7 B= (387-252)/5

OpenStudy (hihi67):

Dude, I got it under control. @AtlasOfAtlantis

OpenStudy (anonymous):

Whoops, I meant divided by nine for A.

OpenStudy (hihi67):

....